OpenStudy (anonymous):

how to find y(s) y''+3y'+2y = 5 y(0)=-1 , y'(0)=2 with laplace

OpenStudy (anonymous):

start by taking the laplace transform of both sides

OpenStudy (anonymous):

L{y'' + 3y' + 2y} = L{5} [L{y}s^2 - sy(0) - y'(0)] + 3[L{y}s - y(0)] + 2[L{y}] = 5/s

OpenStudy (anonymous):

plug values for y(0) and y'(0)

OpenStudy (anonymous):

im stuck in replace initial conditios :S

OpenStudy (anonymous):

Just plug them in after you take the laplace of each side. Then solve the LHS for L{y}

OpenStudy (anonymous):

i have (s^2+3s+2)y(s) + S - 2 - 3 = 5/S correct ?

OpenStudy (anonymous):

I think that it should be +3 on the left not -3

OpenStudy (anonymous):

I have L{y}(s^2+3s+2) = 5/s - 1 - s

OpenStudy (anonymous):

The solution that I came up with is \[y = -5e^{-t}+{3 \over 2}e^{-2t} + {5 \over 2}\]

OpenStudy (anonymous):

thanks